A cell phone is allocated 642 radio frequencies to operate in a city with a hexagon pattern cell design. 42 of the frequencies is reserved for control channel and each phone cell uses 2 frequencies per call ( 1 for transmitting and 1 for receiving). In an advance mobile phone service, use the frequencies distribution to answer the following question. If the distance between centers of adjacent cells is 2km and the minimum distance between centers of cells with the same cochannel is 7km, calculate

a). cell pattern Reuse factor

b). the area of each cell

c). the radius of each cell


1
Expert's answer
2021-10-08T04:35:28-0400

Given,

Allocated frequency (f)= 642 radio frequencies

Reserved frequencies = 42 for control channel

Reserved frequencies for the phone call = 2

1 for transmitting and 1 for receiving.

Distance between the adjacent cell = 2km

The minimum distance between the centers of cells with same co-channel = 7km

a) N=4

total number of frequency available per cell = 660/4 =165channels

b) N=7

total number of frequency available per cell = 660/7 = 95 channel

c) N=12

total number of frequency available per cell = 660/12 =55 channel